OpenStudy (anonymous):

What is the value of 3√ 125 ? A. 5 B. 41.67 C. 375 D. 1,953,125

OpenStudy (luigi0210):

Well \[\LARGE \sqrt{125}=\sqrt{25*5}\] You could find the root of 25.

ganeshie8 (ganeshie8):

if its \(\large \sqrt[3]{125}\) , then convert it to exponent :- \(\large \sqrt[3]{125} = (125)^{\frac{1}{3}}\)